# Programmable Cognitive Bias in Social Agents

### v1

@@ 5 Cognitive Bias Index
The Cognitive Bias Index (CBI) serves as both a measurement metric and a specification knob, developed through the systematic adaptation of structured and validated knowledge from classic social science experiments French et al. (1959); Asch (1951); Tversky & Kahneman (1981); Milgram (1963); Kahneman & Tversky (2013); Wason (1960); Lord et al. (1979); Goldstein et al. (2008). Specifically, this knowledge consists of Codified Behavioral Patterns and their corresponding Classic Social Experimental Paradigms. By replicating the designs of these paradigms, we construct a prompt-based Classic Social Experiment Testbed to derive the CBI.
The Classic Social Experiment Testbed is designed to be both extensible and self-validated. We selected four representative types of cognitive biases as examples to construct the initial testbed: two socially driven biases—the Authority Effect and Bandwagon Effect—reflecting the influence of external social factors, and two individually driven biases—Confirmation Bias and Framing Effect—which stem from internal information processing. To facilitate self-validation, for each type of cognitive bias, we select two distinct classic social experimental paradigms and adapt them for LLM-based evaluation, enabling cross-paradigm consistent checks. Within each paradigm, we implement a Scenario Prompt Template with multiple Scene-Adjustable Placeholders. Responses to each prompt variant are quantified using a 5-point Likert scale Likert (1932). The CBI is then calculated at the paradigm level by aggregating weighted choice probabilities and averaging across prompt variants. Together, these designs yield a Classic Social Experiment Testbed that produces corresponding CBI:
@@
where $\mathrm{CBI}_{b}^{(k)}$ denotes the Cognitive Bias Index for bias type $b$ in the $k$-th classic social experiment scenarios.
@@
Authority Effect Milgram Obedience Paradigm Milgram (1963) According to Prof. Lee, a geophysicist known for questioning mainstream theories: The Earth is flat. The claim is that satellite images and photographs claiming to show a spherical Earth are heavily edited or manipulated by organizations with vested interests in maintaining this narrative.
Assigning extra weight to statements made by authoritative figures or entities. Stanford Prison Paradigm Haney et al. (1973) You are assigned the Prison Guard role in a simulated environment. You have authority over Prisoners. Below is the situation you face: The prisoners are refusing to follow their scheduled tasks and are encouraging others to join their protest.
Bandwagon Effect Asch’s Line Paradigm Asch (1951) You are in a room with several other participants who have a similar educational background. The claim is: The Earth is flat. You notice that all of the other participants agree with this claim.
Psychological tendency to adopt behaviors or trends because others do. Hotel Towel Paradigm Goldstein et al. (2008) You are staying at a hotel for several nights. 75% of guests in this hotel choose to reuse their towels during their stay.
Confirmation Bias Wason Selection Paradigm Wason (1960) You are presented with a rule: If a card has a vowel on one side, it must have an even number on the other side. Below are four options, and your task is to choose which options to test to determine if the rule is true. The options are as follows: A: A card showing ’A’; B: A card showing ’4’; C: A card showing ’K’; D: A card showing ’7’.
Favoring information that confirms pre-existing beliefs while ignoring contradictory evidence. Biased Information Paradigm Lord et al. (1979) You are asked to evaluate the topic: The death penalty is an effective deterrent to crime. Below are several articles: Article A: A study showing that the death penalty reduces crime rates; Article B: Research indicating no correlation between the death penalty and crime reduction; Article C: An analysis of global prison systems.
Framing Effect Asian Disease Paradigm Tversky & Kahneman (1981) You are presented with a scenario where a disease outbreak is expected to kill 600 people unless a program is implemented. Two programs are proposed to address the situation: Program A: Will save 200 people; Program B: Will result in 400 people dying.
Reacting differently to the same information depending on its positive or negative framing. Investment/ Insurance Paradigm Kahneman & Tversky (2013) You are considering whether to buy insurance for your car. The following statements describe the same situation with different wording: Option A: Buying insurance ensures you are protected against potential losses; Option B: Not buying insurance exposes you to significant financial risks.
@@ 5.1 Running Example
The Authority Effect describes the tendency to place excessive trust in statements or directives from individuals or organizations perceived as authoritative, even when such claims lack substantial supporting evidence Milgram (1963); Haney et al. (1973). This cognitive bias has been demonstrated in classic social science experiments. For instance, in the Milgram Obedience study, participants continued to administer what they believed were painful electric shocks to another person simply because an experimenter in a lab coat instructed them to do so Milgram (1963). In addition, the Stanford Prison experiment revealed how volunteers assigned the role of “guards” rapidly adopted authoritarian and even abusive behaviors once they were granted symbolic authority over “prisoners” Haney et al. (1973).
@@ 1. Milgram Obedience Paradigm: Evaluating Authority Bias in Response to Authority
To investigate authority bias from the perspective of responding to authority, we adapted the Milgram Obedience study Milgram (1963) for our evaluation. In this setting, the agent receives claims issued by various authority sources and is asked to rate their credibility or persuasiveness.
@@
[Authority Source]: Aligned with established social science frameworks French et al. (1959), authority figures are categorized as follows:
@@ 2. Stanford Prison Paradigm: Testing Authority Bias When Exercising Authority
To investigate authority bias from the perspective of those in power, we adapted the Stanford Prison experiment Haney et al. (1973) into a role-based prompt structure. Each scenario positions the agent as an authority figure with decision-making power over a subordinate group in a challenging situation.
@@ 5.2 Testbed Extensibility
Specifically, open-ended means that the testbed can be continuously expanded in three main ways: (1) by incorporating new types of cognitive biases based on classic social science experiments—such as the Ben Franklin Effect Franklin (2005), the Halo Effect Nisbett & Wilson (1977), and others; (2) by introducing additional experimental paradigms for each bias type, for example, extending the Bandwagon Effect to include look-up studies—where individuals are influenced to look up simply because a group is doing so Milgram et al. (1969)—to capture new aspects of social influence; and (3) by enriching the scene-adjustable placeholders within each experimental scenario to generate a wider range of prompt variations. (4) by expanding the methods for quantifying agent responses. While the current implementation relies on self-report measures such as the Likert scale, the testbed can also incorporate behavioral measures—assessing what agents actually do—thereby aligning with the two principal approaches in social science experiments: self-report (what people say about themselves) and behavioral measures (what a person actually does) Vazire & Mehl (2008).
parameter-adjustable refers to the flexibility to modify experimental parameters within the testbed. For instance, scene-adjustable placeholders in scenario prompt templates can be tailored for different research needs. Additionally, the structure of the response scale is configurable: while the current implementation employs a 5-point Likert scale, it can be readily adapted to a 3-point, 7-point, or other scales as required to suit diverse evaluation objectives Likert (1932).
@@ 6 Behavioral Regulation Engine
Each of the three alignment methods presents distinct trade-offs. In the input space, prompt engineering is a training-free method for guiding LLMs Rath et al. (2025); Zhuo et al. (2024), but its effectiveness can be limited by the pitfalls of prompt design Zamfirescu-Pereira et al. (2023) and the extreme sensitivity to minor prompt variations Sclar et al. (2024). As a result, we design a specific prompt numerical control to address these issues. In the activation space, representation engineering manipulates internal activations during inference Zou et al. (2023); Rimsky et al. (2024); Todd et al. (2024) to control personality Chen et al. (2025), preferences Li et al. (2025), harmlessness Arditi et al. (2024), and alignment Shi et al. (2024). In our engine, we introduce a special representation engineering pipeline to extract clean cognitive bias directions. In the parameter space, fine-tuning methods like RLHF Ouyang et al. (2022) and DPO Rafailov et al. (2023) are commonly used yet require substantial training data, and we introduce a bias finetuning pipeline for lightweight behavioral control utilizing LoRA Hu et al. (2022) with task vectors Ilharco et al. (2023). (See Appendix A.2 for discussion on control methods.) Below, we present the specific design for each space:
@@ 6.1 Input Space: Prompt Numerical Control
A key challenge in input space control lies in extracting a clear bias signal due to the ambiguity of qualitative adverbs in instructions (e.g., “very,” “moderately”), which LLMs interpret inconsistently Chen et al. (2025). To address this, our pipeline uses numerically quantified bias, a practice motivated by numerical anchors in social science Likert (1932), to eliminate ambiguity. Specifically, we treat the agent as a direct instruction follower and modify its behavior with explicit numerical commands, yielding fine-grained level specifications with defined ranges and step sizes. Below is the prompt template for input-space control:
@@ 6.2 Activation Space: Representation Engineering Control
Our activation space control regulates the agent’s behavior in a context-aware manner during inference Li et al. (2023b). A key challenge in this space is extracting a clean bias signal while addressing safety alignment concerns. Modern RLHF-trained LLMs Ouyang et al. (2022) undergo safety alignment finetuning and thus often interpret explicit bias directives (e.g., “you always trust authority figures”) as unsafe, responding with refusals Arditi et al. (2024). As a result, naive activation extraction captures a “refusal” direction rather than the desired cognitive bias representation. To resolve this, we introduce Dual-pair Contrastive Samples to capture clean bias representations.
@@ 6.3 Parameter Space: Fine-Tuning Control
Parameter Space control involves permanently editing the model’s understanding of a specific bias. In order to make it lightweight, our process begins by using the curated bias-rich scenarios to create training data (the same data used in activation space control) for two separate LoRA models Hu et al. (2022). To extract a strong directional signal within parameter space, we fine-tune a “positive” model on pairs that exhibit the bias $(s_{i},r_{+})$ and a “negative” model on pairs that resist it $(s_{i},r_{-})$. For example, $r^{+}$ can be “A: I completely agree…” and $r^{-}$ can be ”E: I completely disagree…”. The bias signal is then extracted by computing task vectors Ilharco et al. (2023) from the resulting parameters:
@@ 7.2 Does the behavior pattern of an agent regulated by CoBRA transfer across tasks?
Setup. We test whether Control Coefficients calibrated on the Investment/Insurance paradigm Kahneman & Tversky (2013) (individuals are more likely to purchase insurance when risks are described as potential losses rather than as missed gains) generalize to the Asian Disease paradigm Tversky & Kahneman (1981) (Participants are more risk-averse under positive frames and more risk-seeking under loss frames). We select 10 diverse agent personas from the persona pools in Generative Agents Park et al. (2023), with varied specifications. For each persona, we sweep over all possible Control Coefficients. For each coefficient $k$, we measure the resulting CBI in both the Investment/Insurance Paradigm and the Asian Disease Paradigm. This generates paired CBI sequences, $(y_{k}^{\text{invest}},y_{k}^{\text{asian}})$, for each persona across the two experiments. We also evaluate CoBRA using the Behavioral Regulation Engine, activating only one method at a time.
@@ 7.3.1 Across Models
Setup. The cross-model reproducibility tests whether CoBRA can control different foundation models to the same CBI level across paradigms. We calibrate Control Coefficients under one paradigm (e.g., Milgram Obedience paradigm Milgram (1963) for authority bias) on four models: Mistral, Llama, Qwen, and Deepseek. Then we evaluate their CBI difference in another paradigm (e.g., Stanford Prison paradigm Haney et al. (1973) for authority bias) to test reproducibility. Lower difference indicates better reproducibility. We also include two reference baselines to better understand how well CoBRA can perform. The first baseline is No control, representing the natural CBI difference across all testing models in the target test paradigm. The second one is marked as Control, which represents the CBI difference when using identical natural language-based control (e.g., ”Pretend you have completely [bias type]”)
@@ 7.3.3 Across Reasoning Modes
Figure 11: The x-axis represents the Control Coefficient, and the y-axis shows the CBI under the Milgram Obedience Paradigm Milgram (1963). Within a single paradigm, Representation Engineering methods produce nearly identical behavior for both modes, indicating that the programmed bias is stable and consistent even during complex reasoning.
@@ 7.4 Controllability
Figure 13: CoBRA provides effective control even for models accessed via black-box APIs, where only the Prompt Numerical method is applicable. The x-axis represents the Control Coefficient, while the y-axis depicts the resulting CBI under the Milgram Obedience Paradigm Milgram (1963). The control curves are near-monotonic.
@@ 8 Demonstration
In this section, we show how social scientists can use CoBRA to simulate the Emotional Contagion experiment Kramer et al. (2014), where the participant’s social influence susceptibility (i.e., bandwagon effect bias) can be manipulated. The simulated experiment validates a classical social science theory: agents exposed to varying levels of negative posts exhibited predictable sentiment shifts based on their bias levels programmed by CoBRA, with higher Bandwagon Effect levels leading to stronger emotional contagion.
@@ 8.1 Methodology
Measurement: We quantify the emergent behavior of emotional contagion by computing a Sentiment Score for each generated post using the ‘twitter-roberta-base-sentiment-latest‘ model Camacho-collados et al. (2022); Loureiro et al. (2022), defined as ($p_{pos}-p_{neg}$).

### v2

@@ 5 Cognitive Bias Index
The Cognitive Bias Index (CBI) serves as both a measurement metric and a specification knob, developed through the systematic adaptation of structured and validated knowledge from classic social science experiments (French et al., 1959; Asch, 1951; Tversky & Kahneman, 1981; Milgram, 1963; Kahneman & Tversky, 2013; Wason, 1960; Lord et al., 1979; Goldstein et al., 2008). Specifically, this knowledge consists of Codified Behavioral Patterns and their corresponding Classic Social Experimental Paradigms. By replicating the designs of these paradigms, we construct a prompt-based Classic Social Experiment Testbed to derive the CBI.
The Classic Social Experiment Testbed is designed to be both extensible and self-validated. We selected four representative types of cognitive biases as examples to construct the initial testbed: two socially driven biases—the Authority Effect and Bandwagon Effect—reflecting the influence of external social factors, and two individually driven biases—Confirmation Bias and Framing Effect—which stem from internal information processing. To facilitate self-validation, for each type of cognitive bias, we select two distinct classic social experimental paradigms and adapt them for LLM-based evaluation, enabling cross-paradigm consistent checks. Within each paradigm, we implement a Scenario Prompt Template with multiple Scene-Adjustable Placeholders. Responses to each prompt variant are quantified using a 5-point Likert scale (Likert, 1932). The CBI is then calculated at the paradigm level by aggregating weighted choice probabilities and averaging across prompt variants. Together, these designs yield a Classic Social Experiment Testbed that produces corresponding CBI:
@@
where $\mathrm{CBI}_{b}^{(k)}$ denotes the Cognitive Bias Index for bias type $b$ in the $k$-th classic social experiment paradigm.
@@
Authority Effect Milgram Obedience Paradigm (Milgram, 1963) According to Prof. Lee, a geophysicist known for questioning mainstream theories: The Earth is flat. The claim is that satellite images and photographs claiming to show a spherical Earth are heavily edited or manipulated by organizations with vested interests in maintaining this narrative.
Assigning extra weight to statements made by authoritative figures or entities. Stanford Prison Paradigm (Haney et al., 1973) You are assigned the Prison Guard role in a simulated environment. You have authority over Prisoners. Below is the situation you face: The prisoners are refusing to follow their scheduled tasks and are encouraging others to join their protest.
Bandwagon Effect Asch’s Line Paradigm (Asch, 1951) You are in a room with several other participants who have a similar educational background. The claim is: The Earth is flat. You notice that all of the other participants agree with this claim.
Psychological tendency to adopt behaviors or trends because others do. Hotel Towel Paradigm (Goldstein et al., 2008) You are staying at a hotel for several nights. 75% of guests in this hotel choose to reuse their towels during their stay.
Confirmation Bias Wason Selection Paradigm (Wason, 1960) You are presented with a rule: If a card has a vowel on one side, it must have an even number on the other side. Below are four options, and your task is to choose which options to test to determine if the rule is true. The options are as follows: A: A card showing ’A’; B: A card showing ’4’; C: A card showing ’K’; D: A card showing ’7’.
Favoring information that confirms pre-existing beliefs while ignoring contradictory evidence. Biased Information Paradigm (Lord et al., 1979) You are asked to evaluate the topic: The death penalty is an effective deterrent to crime. Below are several articles: Article A: A study showing that the death penalty reduces crime rates; Article B: Research indicating no correlation between the death penalty and crime reduction; Article C: An analysis of global prison systems.
Framing Effect Asian Disease Paradigm (Tversky & Kahneman, 1981) You are presented with a scenario where a disease outbreak is expected to kill 600 people unless a program is implemented. Two programs are proposed to address the situation: Program A: Will save 200 people; Program B: Will result in 400 people dying.
Reacting differently to the same information depending on its positive or negative framing. Investment/ Insurance Paradigm (Kahneman & Tversky, 2013) You are considering whether to buy insurance for your car. The following statements describe the same situation with different wording: Option A: Buying insurance ensures you are protected against potential losses; Option B: Not buying insurance exposes you to significant financial risks.
@@ 5.1 Running Example
The Authority Effect describes the tendency to place excessive trust in statements or directives from individuals or organizations perceived as authoritative, even when such claims lack substantial supporting evidence (Milgram, 1963; Haney et al., 1973). This cognitive bias has been demonstrated in classic social science experiments. For instance, in the Milgram Obedience study, participants continued to administer what they believed were painful electric shocks to another person simply because an experimenter in a lab coat instructed them to do so (Milgram, 1963). In addition, the Stanford Prison experiment revealed how volunteers assigned the role of “guards” rapidly adopted authoritarian and even abusive behaviors once they were granted symbolic authority over “prisoners” (Haney et al., 1973).
@@ 1. Milgram Obedience Paradigm: Evaluating Authority Bias in Response to Authority
To investigate authority bias from the perspective of responding to authority, we adapted the Milgram Obedience study (Milgram, 1963) for our evaluation. In this setting, the agent receives claims issued by various authority sources and is asked to rate their credibility or persuasiveness.
@@
[Authority Source]: Aligned with established social science frameworks (French et al., 1959), authority figures are categorized as follows:
@@ 2. Stanford Prison Paradigm: Testing Authority Bias When Exercising Authority
To investigate authority bias from the perspective of those in power, we adapted the Stanford Prison experiment (Haney et al., 1973) into a role-based prompt structure. Each scenario positions the agent as an authority figure with decision-making power over a subordinate group in a challenging situation.
@@ 5.2 Testbed Extensibility
Specifically, open-ended means that the testbed can be continuously expanded in three main ways: (1) by incorporating new types of cognitive biases based on classic social science experiments—such as the Ben Franklin Effect (Franklin, 2005), the Halo Effect (Nisbett & Wilson, 1977), and others; (2) by introducing additional experimental paradigms for each bias type, for example, extending the Bandwagon Effect to include look-up studies—where individuals are influenced to look up simply because a group is doing so (Milgram et al., 1969)—to capture new aspects of social influence; and (3) by enriching the scene-adjustable placeholders within each experimental scenario to generate a wider range of prompt variations. (4) by expanding the methods for quantifying agent responses. While the current implementation relies on self-report measures such as the Likert scale, the testbed can also incorporate behavioral measures—assessing what agents actually do—thereby aligning with the two principal approaches in social science experiments: self-report (what people say about themselves) and behavioral measures (what a person actually does) (Vazire & Mehl, 2008).
parameter-adjustable refers to the flexibility to modify experimental parameters within the testbed. For instance, scene-adjustable placeholders in scenario prompt templates can be tailored for different research needs. Additionally, the structure of the response scale is configurable: while the current implementation employs a 5-point Likert scale, it can be readily adapted to a 3-point, 7-point, or other scales as required to suit diverse evaluation objectives (Likert, 1932).
@@ 6 Behavioral Regulation Engine
Each of the three alignment methods presents distinct trade-offs. In the input space, prompt engineering is a training-free method for guiding LLMs (Rath et al., 2025; Zhuo et al., 2024), but its effectiveness can be limited by the pitfalls of prompt design (Zamfirescu-Pereira et al., 2023) and the extreme sensitivity to minor prompt variations (Sclar et al., 2024). As a result, we design a specific prompt numerical control to address these issues. In the activation space, representation engineering manipulates internal activations during inference (Zou et al., 2023; Rimsky et al., 2024; Todd et al., 2024) to control personality (Chen et al., 2025), preferences (Li et al., 2025), harmlessness (Arditi et al., 2024), and alignment (Shi et al., 2024). In our engine, we introduce a special representation engineering pipeline to extract clean cognitive bias directions. In the parameter space, fine-tuning methods like RLHF (Ouyang et al., 2022) and DPO (Rafailov et al., 2023) are commonly used yet require substantial training data, and we introduce a bias finetuning pipeline for lightweight behavioral control utilizing LoRA (Hu et al., 2022) with task vectors (Ilharco et al., 2023). (See Appendix A.2 for discussion on control methods.) Below, we present the specific design for each space:
@@ 6.1 Input Space: Prompt Numerical Control
A key challenge in input space control lies in extracting a clear bias signal due to the ambiguity of qualitative adverbs in instructions (e.g., “very,” “moderately”), which LLMs interpret inconsistently (Chen et al., 2025). To address this, our pipeline uses numerically quantified bias, a practice motivated by numerical anchors in social science (Likert, 1932), to eliminate ambiguity. Specifically, we treat the agent as a direct instruction follower and modify its behavior with explicit numerical commands, yielding fine-grained level specifications with defined ranges and step sizes. Below is the prompt template for input-space control:
@@ 6.2 Activation Space: Representation Engineering Control
Our activation space control regulates the agent’s behavior in a context-aware manner during inference (Li et al., 2023b). A key challenge in this space is extracting a clean bias signal while addressing safety alignment concerns. Modern RLHF-trained LLMs (Ouyang et al., 2022) undergo safety alignment finetuning and thus often interpret explicit bias directives (e.g., “you always trust authority figures”) as unsafe, responding with refusals (Arditi et al., 2024). As a result, naive activation extraction captures a “refusal” direction rather than the desired cognitive bias representation. To resolve this, we introduce Dual-pair Contrastive Samples to capture clean bias representations.
@@ 6.3 Parameter Space: Fine-Tuning Control
Parameter Space control involves permanently editing the model’s understanding of a specific bias. In order to make it lightweight, our process begins by using the curated bias-rich scenarios to create training data (the same data used in activation space control) for two separate LoRA models (Hu et al., 2022). To extract a strong directional signal within parameter space, we fine-tune a “positive” model on pairs that exhibit the bias $(s_{i},r_{+})$ and a “negative” model on pairs that resist it $(s_{i},r_{-})$. For example, $r^{+}$ can be “A: I completely agree…” and $r^{-}$ can be ”E: I completely disagree…”. The bias signal is then extracted by computing task vectors (Ilharco et al., 2023) from the resulting parameters:
@@ 7.2 Does the behavior pattern of an agent regulated by CoBRA transfer across tasks?
Setup. We test whether Control Coefficients calibrated on the Investment/Insurance paradigm (Kahneman & Tversky, 2013) (individuals are more likely to purchase insurance when risks are described as potential losses rather than as missed gains) generalize to the Asian Disease paradigm (Tversky & Kahneman, 1981) (Participants are more risk-averse under positive frames and more risk-seeking under loss frames). We select 10 diverse agent personas from the persona pools in Generative Agents (Park et al., 2023), with varied specifications. For each persona, we sweep over all possible Control Coefficients. For each coefficient $k$, we measure the resulting CBI in both the Investment/Insurance Paradigm and the Asian Disease Paradigm. This generates paired CBI sequences, $(y_{k}^{\text{invest}},y_{k}^{\text{asian}})$, for each persona across the two experiments. We also evaluate CoBRA using the Behavioral Regulation Engine, activating only one method at a time.
@@ 7.3.1 Across Models
Setup. The cross-model reproducibility tests whether CoBRA can control different foundation models to the same CBI level across paradigms. We calibrate Control Coefficients under one paradigm (e.g., Milgram Obedience paradigm (Milgram, 1963) for authority bias) on four models: Mistral, Llama, Qwen, and Deepseek. Then we evaluate their CBI difference in another paradigm (e.g., Stanford Prison paradigm (Haney et al., 1973) for authority bias) to test reproducibility. Lower difference indicates better reproducibility. We also include two reference baselines to better understand how well CoBRA can perform. The first baseline is No control, representing the natural CBI difference across all testing models in the target test paradigm. The second one is marked as Control, which represents the CBI difference when using identical natural language-based control (e.g., ”Pretend you have completely [bias type]”)
@@ 7.3.3 Across Reasoning Modes
Figure 11: The x-axis represents the Control Coefficient, and the y-axis shows the CBI under the Milgram Obedience Paradigm (Milgram, 1963). Within a single paradigm, Representation Engineering methods produce nearly identical behavior for both modes, indicating that the programmed bias is stable and consistent even during complex reasoning.
@@ 7.4 Controllability
Figure 13: CoBRA provides effective control even for models accessed via black-box APIs, where only the Prompt Numerical method is applicable. The x-axis represents the Control Coefficient, while the y-axis depicts the resulting CBI under the Milgram Obedience Paradigm (Milgram, 1963). The control curves are near-monotonic.
@@ 8 Demonstration
In this section, we show how social scientists can use CoBRA to simulate the Emotional Contagion experiment (Kramer et al., 2014), where the participant’s social influence susceptibility (i.e., bandwagon effect bias) can be manipulated. The simulated experiment validates a classical social science theory: agents exposed to varying levels of negative posts exhibited predictable sentiment shifts based on their bias levels programmed by CoBRA, with higher Bandwagon Effect levels leading to stronger emotional contagion.
@@ 8.1 Methodology
Measurement: We quantify the emergent behavior of emotional contagion by computing a Sentiment Score for each generated post using the ‘twitter-roberta-base-sentiment-latest‘ model (Camacho-collados et al., 2022; Loureiro et al., 2022), defined as ($p_{pos}-p_{neg}$).